OpenStudy (brinazarski):

Computer the smallest value of n for which a regular polygon of n sides has at least 2006 diagonals. I got 2006, is this correct?

OpenStudy (brinazarski):

*compute

jimthompson5910 (jim_thompson5910):

let d = number of diagonals and n = number of sides

jimthompson5910 (jim_thompson5910):

the formula is d = n(n-3)/2 so because d = 2006, we know that d = n(n-3)/2 2006 = n(n-3)/2 2006*2 = n^2 - 3n 4012 = n^2 - 3n n^2 - 3n - 4012 = 0 Solve that for n to get your answer

OpenStudy (brinazarski):

3/2 + \[\sqrt{4014.25}\]

OpenStudy (brinazarski):

?

jimthompson5910 (jim_thompson5910):

use a calculator to find the decimal form of that number

OpenStudy (brinazarski):

I did, it seems right...

jimthompson5910 (jim_thompson5910):

you should get n = 64.85810919 and this rounds up to n = 65

OpenStudy (brinazarski):

yup.... but why round up? It doesn't say round up

jimthompson5910 (jim_thompson5910):

because you can't have a decimal/fractional number of sides

OpenStudy (brinazarski):

Why not?

jimthompson5910 (jim_thompson5910):

how do you have 0.85 of a side?

OpenStudy (brinazarski):

By have .85 of a side. You can use a ruler and make a square with each side 1.5 inches, can't you?

jimthompson5910 (jim_thompson5910):

no, n = number of sides, not length of the sides

OpenStudy (brinazarski):

Oh... that's true. I'm sorry. Thank you! :)

jimthompson5910 (jim_thompson5910):

no worries, yw